please help me:
Find the two numbers whose sum is 10; if twice the first number is 3 less than the other number

Answer by fractalier(6550)

Call the numbers x and y. Then we have
x + y = 10 (or y = 10-x)
and
2x = y - 3
Now substitute and get
2x = 10-x - 3
2x = 7 - x
3x = 7
x = 7/3
y = 10 - x = 30/3 - 7/3 = 23/3